OpenStudy (anonymous):

If f (x) = |x-5| + 3, find f(2) (Points 1) A. 11 B. 6 C. 4 D. 0 PLEASE HELP I HAVE LIMITED TIME TO GET THIS DONE.

OpenStudy (hockeychick23):

2-5= -3 but its in an absolute value sign so its 3+3=6 or B

OpenStudy (anonymous):

\[f(2)=\left| 2-5 \right|+3\]\[=\left| -3 \right|+3\]\[=3+3\]\[=?\]

OpenStudy (anonymous):

solve it @bvblover83014

OpenStudy (butterflydreamer):